PART A A sample of ideal gas is in a sealed container. The pressure of the...

PART A

A sample of ideal gas is in a sealed container. The pressure of the gas is 445 torr , and the temperature is 26 ∘C . If the temperature changes to 98 ∘C with no change in volume or amount of gas, what is the new pressure, P2, of the gas inside the container?

Express your answer with the appropriate units.

PART B

Using the same sample of gas (P1 = 445 torr , T1 = 26 ∘C ), we wish to change the pressure to 4450 torr with no accompanying change in volume or amount of gas. What temperature T2, in Celsius, is needed to reach this pressure?

Express your answer with the appropriate units.
Answer #1

A)

Given:

Pi = 445 torr

Ti = 26.0 oC

= (26.0+273) K

= 299 K

Tf = 98.0 oC

= (98.0+273) K

= 371 K

use:

Pi/Ti = Pf/Tf

445 torr / 299 K = Pf / 371 K

Pf = 552 torr

Answer: 552 torr

B)

Given:

Ti = 26.0 oC

= (26.0+273) K

= 299 K

Pi = 445 torr

Pf = 4450 torr

use:

Pi/Ti = Pf/Tf

445.0 torr / 299.0 K = 4450.0 torr / Tf

Tf = 2990 K

= (2990-273) oC

= 2717 oC

Answer: 2717 oC